# Fully Coupled Implicit Hydro-Mechanical Multiphase Flow Simulation in Deformable Porous Media Using DEM

### v1

@@ Abstract
Knowledge of the underlying mechanisms of multiphase flow dynamics in porous media is crucial for optimizing subsurface engineering applications like geological carbon sequestration. However, studying the micro-mechanisms of multiphase fluid–grain interactions in the laboratory is challenging due to the difficulty in obtaining mechanical data such as force and displacement. Transitional discrete element method models coupled with pore networks offer insights into these interactions but struggle with accurate pressure prediction during pore expansion from fracturing and efficient simulation during the slow drainage of compressible fluids. To address these limitations, we develop an advanced two-way coupled hydro-mechanical discrete element method model that accurately and efficiently captures fluid–fluid and fluid–grain interactions in deformable porous media. Our model integrates an unconditionally stable implicit finite volume approach, enabling significant timesteps for advancing fluids. A pressure-volume iteration scheme dynamically balances injection-induced pressure buildup with substantial pore structure deformation, while flow front-advancing criteria precisely locate the fluid–fluid interface and adaptively refine timesteps, particularly when capillary effects block potential flow paths. The model is validated against benchmark Hele-Shaw experiments in both rigid and deformable porous media, providing quantitative insights into the micro-mechanisms governing multiphase flow. For the first time, grain-scale inputs such as viscous and capillary pressures, energies, contact forces, and flow resistances are utilized to provide a detailed understanding of micro-scale fluid–fluid and fluid–grain flow patterns and their transitions.
@@ 2 Methodology and formulations
Gravity and inertia are neglected for the lateral fluid invasion in Hele-Shaw tests.
@@ 3 Numerical implementation
Figure 2 outlines the complete numerical implementation of the coupled HM-DEM model. We begin by discretizing the porous media into a grain-contact interconnected pore network in PFC2D. The implicit pressure solver then provides pressure solutions for moving grains and advancing fluids. Concurrently, the grain motion solver manages the dynamic interactions between grains and contacts in a time-stepping manner, repeatedly applying the laws of motion and force-displacement until force equilibrium is achieved. However, the fluid-driven re-arrangement of the pore structure continually challenges pressure solution accuracy. We introduce the pressure-volume iteration scheme to synchronize the fluid pressure and grain motion solvers to address this. The flow front-advancing criteria enable precise location of the fluid–fluid interface in deformable pore media. A timestep adjustment occurs when the invading fluid simultaneously advances into two continuous domains or when the capillary effect blocks all potential flow paths of a specific domain. Before initiating the next flow timestep, our model updates the pore volume of each domain (), the throat conductance of each pipe (), and the fluid bulk modulus () and viscosity () of the front domains. The simulation concludes once the flow front reaches the outflow boundary.
@@ 3.1 Implicit pressure solver
where is the set of surrounding domain indices around domain , is the set of domain indices connected to domain by a pipe containing fluid–fluid interface, is the subset of where flow is blocked by capillary effects, and is the set of inflow domain indices.
@@ 4 Numerical simulations
In this section, we construct both linear and radial flow models to validate the developed HM-DEM framework and explore the underlying mechanisms of multiphase flow in porous media. The numerical simulations involve two distinct yet complementary schemes: (i) linear injection tests in fixed-grain porous media to verify the feasibility of producing realistic fluid–fluid displacement patterns and (ii) radial injection tests in movable-grain porous media to demonstrate the capability to replicate complex fluid–grain interactions.
@@
All simulations are conducted on a computer with an 8-Core i7-4790 CPU and 32 GB of RAM. Injection flow simulations take several hours to a full day, depending on the sample dimension, selected timestep, and infiltration scale of the reproduced flow patterns.
@@ 4.1.1 Model setup
To reproduce the observations of Lenormand’s experiment—namely, viscous fingering, capillary fingering, and stable displacement—we conduct two drainage scenarios with viscosity ratios of and 101.8. Each scenario includes two injection tests with different modified capillary numbers: (i) air displacing very-viscous oil with of and , and (ii) glucose solution (mainly water) displacing less-viscous oil with of 0.25 and . Table 1 summarizes the flow properties for these drainage cases, ensuring that all hydraulic properties are consistent with Lenormand’s experiment. The fluid bulk modulus values are drawn from the literature [69, 70].
@@ 4.1.2 Results
Figure 6 and Figure 9 offer quantitative insights into the fundamental mechanisms driving different flow patterns and their transitions in Case 1 and Case 2, respectively. They depict the temporal evolution of the average injection pressure across all inlet domains (), highlighting three specific time instances ( 0.1, 0.5, and 0.9) to illustrate how the flow pattern and pressure field evolve over time.
@@
To explore the micro-mechanisms behind viscous fingering (Figure 7) and capillary fingering (Figure 8), we focus on the evolution of the fingertip during the second time instance () shown in Figure 6. In Figure 7a, the potential invasion paths at the viscous fingertip are multi-directional. As the invading fluid displaces the defending fluid into adjacent domains, viscous pressure increases. However, the rate of pressure dissipation is primarily influenced by the discharge distance between the defending-fluid domain and the outflow boundary. Shorter distances offer less resistance to dissipation, causing the viscous pressure in the defending-fluid domain at the fingertip front to dissipate more rapidly. As a result, the front domains develop significant pressure differences, leading to active flow paths, while the domains behind them remain with nearly zero pressure difference (Figure 7b). Over time, the domain exhibiting a relatively significant pressure difference ( kPa) and aperture width ( m) fills first (Figure 7c). The flow then infiltrates the next preferential domain ( kPa and m) (Figure 7d), which also presents a larger pressure difference due to its short dissipation distance to the outflow boundary. This domain fills shortly afterward (Figure 7e). Given the randomly distributed aperture widths, preferential flow paths are likely to form in vertical directions, potentially developing sub-branches if the aperture widths of the front domain are small. In Figure 7f, the side pipe with kPa and m dominates. Nevertheless, the main branch of viscous fingering continues to advance, as significant pressure differences are concentrated at the leading fingertip.
@@ 4.2.1 Model setup
We select two testing groups from Huang’s experiment—Group A and Group B—with individual invading fluid viscosities of 5 and 176 cP, corresponding to viscosity ratios () of 277.8 and 9777.8, respectively. Each group is simulated with two experimental injection velocities of 8.22 and 165.80 mm/s, corresponding to two modified capillary numbers: (i) A1 and A3 with of 0.04 and 0.89, and (ii) B1 and B3 with of 1.56 and 31.47. These simulations cover typical fluid–grain interaction regimes observed in Huang’s experiment: (i) pure fluid infiltration with no fracturing, (ii) infiltration-dominated with fracturing, and (iii) fracturing-dominated, guiding preferential invasion paths.
@@ 4.2.2 Model calibration
For hydraulic properties listed in Table 3, in addition to properties defined in Huang’s experiment, we calibrate the aperture width-related micro-parameters (, , and ) by conducting a series of sample compactions and Darcy flow tests under varying confining stresses from 0.1 to 2.1 MPa. First, we use the servo mechanism to isotopically compact the initial assembly under specific confining stress and calculate the Kozeny-Carman permeability using the simulated porosity via Eq. (20). Next, we perform Darcy flow tests on the compacted assembly and predict the Darcy-flow permeability using Eq. (21). Since the effect of hydraulic constants on permeability varies with the number and magnitude of compressive contacts, we arbitrarily test different combinations until finalizing a set of hydraulic constants that results in all Darcy-flow permeabilities matching the corresponding Kozeny-Carman values (with an error within 1), as shown in Figure 12b. In summary, across varying confining stress levels, our DEM-generated sample demonstrates consistent fluid conductance with the experimental sample, indicating similar fluid–grain interactions. Following Huang’s experiment, we select a confining stress of 0.14 MPa to prepare the sample for injection simulations, where the predicted Darcy-flow permeability ( m2) matches the experiment value.
@@ 4.2.3 Results
Table 4 lists the fluid flow timestep and total simulation time of movable-grain tests across four distinct combinations of injection velocity and invading fluid viscosity, highlighting our modeling efficiency. The timestep for each velocity ensures that our model performs at least three fluid-grain interaction steps to fill one domain. Compared with other HM-DEM coupled simulations of Huang’s experiment, as seen in Zhang et al. [51], our model not only selects more realistic fluid compressibility but also allows for a timestep more than 80 times larger by employing the implicit pressure solver, significantly enhancing simulation efficiency. Besides, in their one-way coupling model, the injection pressure under large (i.e. B3) causes significant volume changes, leading to pressure fluctuations (sudden drops). Our two-way pressure-volume iteration scheme avoids this issue, producing a smoother pressure development curve.
| Test | Velocity [mm/s] | [cP] | Flow timestep [s] | Simulation time [s] | |
| A1 | 8.22 | 5 | 0.04 | ||
| A3 | 165.80 | 5 | 0.89 | ||
| B1 | 8.22 | 176 | 1.56 | ||
| B3 | 165.80 | 176 | 31.47 |
@@ 5 Conclusions
Moreover, our model offers quantitative insights into the underlying micro-mechanisms governing multiphase flow in porous media, marking the first time that micro-scale interactions have been visually reviewed in terms of viscous and capillary pressures, energies, contact forces, and flow resistances. In simulating Lenormand’s experiment, we identify pattern-transition signals from viscous fingering and stable displacement to capillary fingering by comparing throat-scale viscous and capillary entry pressures. In simulating Huang’s experiment, we elucidate the governing regimes of fluid–grain interactions by comparing pore-scale residual energies and contact forces for fracturing, and flow resistances for fluid dissipation. Notably, while high flow rates and increased invading fluid viscosity in deformable porous media enhance fracturing and influence fluid infiltration paths, this effect contrasts with the positive impact on production efficiency observed in rigid porous media. This discrepancy underscores the need for further systematic parametric studies using our model to fully understand the balance between operational efficiency and system stability in geological processes.
@@ Acknowledgments
This work was supported by the HKU Postgraduate Scholarship. Kang Duan would like to appreciate the support of the National Key RD Program of China (No. 2023YFB2390300). We acknowledge the helpful discussions with Dr. Zhibing Yang of Wuhan University.

### v2

@@ Abstract
Knowledge of the underlying mechanisms of multiphase flow dynamics in porous media is crucial for optimizing subsurface engineering applications like geological carbon sequestration. However, studying the micro-mechanisms of multiphase fluid–grain interactions in the laboratory is challenging due to the difficulty in obtaining mechanical data such as force and displacement. Traditional discrete element method models coupled with pore networks offer insights into these interactions but struggle with accurate pressure prediction during pore expansion from fracturing and efficient simulation during the slow drainage of compressible fluids. To address these limitations, we develop an advanced two-way coupled hydro-mechanical discrete element method model that accurately and efficiently captures fluid–fluid and fluid–grain interactions in deformable porous media. Our model integrates an unconditionally stable implicit finite volume approach, enabling significant timesteps for advancing fluids. A pressure-volume iteration scheme dynamically balances injection-induced pressure buildup with substantial pore structure deformation, while flow front-advancing criteria precisely locate the fluid–fluid interface and adaptively refine timesteps, particularly when capillary effects block potential flow paths. The model is validated against benchmark Hele-Shaw experiments in both rigid and deformable porous media, providing quantitative insights into the micro-mechanisms governing multiphase flow. For the first time, grain-scale inputs such as viscous and capillary pressures, energies, contact forces, and flow resistances are utilized to provide a detailed understanding of micro-scale fluid–fluid and fluid–grain flow patterns and their transitions.
@@ 2 Methodology and formulations
Gravity and inertia are neglected for the lateral fluid invasion in Hele-Shaw tests. Gravity has minimal impact in the lateral direction, and inertia is negligible due to the low Reynolds numbers (), where viscous forces dominate [33]. For scenarios such as field-scale CO2 plume migration (upward) [14] or high inflow rate Hele-Shaw tests () [28], where gravity or inertia significantly influence flow behavior, the model would need to be extended, though this is beyond the scope of the current study.
@@ 3 Numerical implementation
Figure 2 outlines the complete numerical implementation of the coupled HM-DEM model. We begin by discretizing the porous media into a grain-contact interconnected pore network in PFC2D. The implicit pressure solver then provides pressure solutions for moving grains and advancing fluids. Concurrently, the grain motion solver manages the dynamic interactions between grains and contacts in a time-stepping manner, repeatedly applying the laws of motion and force-displacement until force equilibrium is achieved. However, the fluid-driven re-arrangement of the pore structure continually challenges pressure solution accuracy. We introduce the pressure-volume iteration scheme to synchronize the fluid pressure and grain motion solvers to address this. The flow front-advancing criteria enable precise location of the fluid–fluid interface in deformable porous media. A timestep adjustment occurs when the invading fluid simultaneously advances into two continuous domains or when the capillary effect blocks all potential flow paths of a specific domain. Before initiating the next flow timestep, our model updates the pore volume of each domain (), the throat conductance of each pipe (), and the fluid bulk modulus () and viscosity () of the front domains. The simulation concludes once the flow front reaches the outflow boundary.
@@ 3.1 Implicit pressure solver
where is the set of surrounding domain indices around domain , is the set of domain indices connected to domain by a pipe containing fluid–fluid interface, is the subset of where flow is blocked by capillary effects, and is the set of inflow domain indices. For simplicity, only the viscous pressure terms include the superscripts and to distinguish values from the current and advanced timesteps during pressure calculation. All other input parameters, such as , , , and , are assumed to represent values at the current timestep unless explicitly stated otherwise.
@@ 4 Numerical simulations
In this section, we construct both linear and radial flow models to validate the developed HM-DEM framework and explore the underlying mechanisms of multiphase flow in porous media. The simulations involve two distinct yet complementary schemes: (i) linear injection tests in fixed-grain porous media to verify the feasibility of producing realistic fluid–fluid displacement patterns and (ii) radial injection tests in movable-grain porous media to demonstrate the capability to replicate complex fluid–grain interactions.
@@
Notably, our numerical simulations are conducted in a two-dimensional framework, while the experiments are performed in three dimensions. In our 2D model, grains are represented as connected disks with artificial throats for fluid migration. This simplification does not fully capture the vertical heterogeneity of pore structures and the complex grain motions present in 3D systems; however, the small thickness-to-length ratio of Hele-Shaw cell—remaining below 0.01 [22]—ensures that the 3D flow behavior remains comparable to the 2D model. By matching permeability and using the same dimensionless parameters ( and ) as in the experiments, we can reproduce the relevant fluid-fluid and fluid-grain displacement patterns.
All simulations are conducted on a computer with an 8-Core i7-4790 CPU and 32 GB of RAM. Injection flow simulations take several hours to a full day, depending on the sample dimension, selected timestep, and infiltration scale of the reproduced flow patterns. The robustness and accuracy of the model are validated through supplementary tests, including a resolution convergence study (A), timestep sensitivity analysis (B), pressure solver performance evaluation (C), and verification using simple flow scenarios, such as single-phase Darcy flow and two-phase flow in regular packings (D). Comprehensive details of these validations are provided in the appendices.
@@ 4.1.1 Model setup
To reproduce the observations of Lenormand’s experiment—namely, viscous fingering, capillary fingering, and stable displacement—we conduct two drainage scenarios with viscosity ratios of and 101.8. Each scenario includes two injection tests with different modified capillary numbers: (i) air displacing very-viscous oil with of and , and (ii) glucose solution (mainly water) displacing less-viscous oil with of 38.6 and . Table 1 summarizes the flow properties for these drainage cases, ensuring that all hydraulic properties are consistent with Lenormand’s experiment. The fluid bulk modulus values are drawn from the literature [69, 70].
@@ 4.1.2 Results
Beyond qualitative morphology, the flow regimes can be quantitatively classified by their fractal dimension using the box-counting method [71] (see E). As increases, the fractal dimension decreases in the low case (Case 1), transitioning from less-branched capillary fingering (1.83) to more-branched viscous fingering (1.63). In contrast, in the high case (Case 2), the fractal dimension increases with , evolving from less-branched capillary fingering (1.84) to flat-front stable displacement (1.96). These values align with experimentally determined ranges: viscous fingering [1.62, 1.65] [72, 73], capillary fingering [1.80, 1.83] [74], and the typical value of 1.95 for stable displacement [75]. This demonstrates that our model successfully captures key regime transitions, including the shift from viscous fingering to capillary fingering and the transition from stable displacement to capillary fingering.
Figure 6 and Figure 9 illustrate the fundamental mechanisms driving different flow patterns and their transitions in Case 1 and Case 2, respectively. They depict the temporal evolution of the average injection pressure across all inlet domains (), highlighting three specific time instances ( 0.1, 0.5, and 0.9) to illustrate how the flow pattern and pressure field evolve over time.
@@
To explore the micro-mechanisms behind viscous fingering (Figure 7) and capillary fingering (Figure 8), we focus on the evolution of the fingertip during the second time instance () shown in Figure 6. This analysis includes grain-scale viscous pressure differences, aperture widths, and capillary entry pressures.
In Figure 7a, the potential invasion paths at the viscous fingertip are multi-directional. As the invading fluid displaces the defending fluid into adjacent domains, viscous pressure increases. However, the rate of pressure dissipation is primarily influenced by the discharge distance between the defending-fluid domain and the outflow boundary. Shorter distances offer less resistance to dissipation, causing the viscous pressure in the defending-fluid domain at the fingertip front to dissipate more rapidly. As a result, the front domains develop significant pressure differences, leading to active flow paths, while the domains behind them remain with nearly zero pressure difference (Figure 7b). Over time, the domain exhibiting a relatively significant pressure difference ( kPa) and aperture width ( m) fills first (Figure 7c). The flow then infiltrates the next preferential domain ( kPa and m) (Figure 7d), which also presents a larger pressure difference due to its short dissipation distance to the outflow boundary. This domain fills shortly afterward (Figure 7e). Given the randomly distributed aperture widths, preferential flow paths are likely to form in vertical directions, potentially developing sub-branches if the aperture widths of the front domain are small. In Figure 7f, the side pipe with kPa and m dominates. Nevertheless, the main branch of viscous fingering continues to advance, as significant pressure differences are concentrated at the leading fingertip.
@@ 4.2.1 Model setup
We select two testing groups from Huang’s experiment—Group A and Group B—with individual invading fluid viscosities of 5 and 176 cP, corresponding to viscosity ratios () of 277.8 and 9777.8, respectively. Each group is simulated with two experimental injection velocities of 8.22 and 165.80 mm/s, corresponding to two modified capillary numbers: (i) A1 and A3 with of 1.6 and 32.3, and (ii) B1 and B3 with of 56.3 and 1138.1. These simulations cover typical fluid–grain interaction regimes observed in Huang’s experiment: (i) pure fluid infiltration with no fracturing, (ii) infiltration-dominated with fracturing, and (iii) fracturing-dominated, guiding preferential invasion paths.
@@ 4.2.2 Model calibration
For hydraulic properties listed in Table 3, in addition to properties defined in Huang’s experiment, we calibrate the aperture width-related micro-parameters (, , and ) by conducting a series of sample compactions and Darcy flow tests under varying confining stresses from 0.1 to 2.1 MPa. First, we use the servo mechanism, a feedback control system that dynamically adjusts boundary positions, to isotopically compact the initial assembly under specific confining stress and calculate the Kozeny-Carman permeability using the simulated porosity via Eq. (20). Next, we perform Darcy flow tests on the compacted assembly and predict the Darcy-flow permeability using Eq. (21). Since the effect of hydraulic constants on permeability varies with the number and magnitude of compressive contacts, we arbitrarily test different combinations until finalizing a set of hydraulic constants that results in all Darcy-flow permeabilities matching the corresponding Kozeny-Carman values (with an error within 1), as shown in Figure 12b. In summary, across varying confining stress levels, our DEM-generated sample demonstrates consistent fluid conductance with the experimental sample, indicating similar fluid–grain interactions. Following Huang’s experiment, we select a confining stress of 0.14 MPa to prepare the sample for injection simulations, where the predicted Darcy-flow permeability ( m2) matches the experiment value.
@@ 4.2.3 Results
Table 4 presents the fluid flow timestep and total simulation time for movable-grain tests across four distinct combinations of viscosity ratio () and modified capillary number (), highlighting our modeling efficiency achieved with reasonable computational time and memory usage. The timestep for each ensures that our model performs at least three fluid-grain interaction steps to fill one domain. Compared with other HM-DEM coupled simulations of Huang’s experiment, as seen in Zhang et al. [51], our model not only selects more realistic fluid compressibility but also allows for a timestep more than 80 times larger by employing the implicit pressure solver, significantly enhancing simulation efficiency. Besides, in their one-way coupling model, the injection pressure under large (i.e. B3) causes significant volume changes, leading to pressure fluctuations (sudden drops). Our two-way pressure-volume iteration scheme avoids this issue, producing a smoother pressure development curve.
| Test | [s] | [s] | [hour] | Memory [MB] | ||
| A1 | 277.8 | 1.6 | 21.5 | 270 to 1200 | ||
| A3 | 277.8 | 32.3 | 25.6 | 270 to 1200 | ||
| B1 | 9777.8 | 56.3 | 16.2 | 270 to 1200 | ||
| B3 | 9777.8 | 1138.1 | 2.0 | 270 to 400 |
@@
Quantitative analysis using the box-counting fractal dimension (see E) shows good agreement with experimental data. The fractal dimension of the circular infiltration zones in A1, A3, and B1 matches the experimental value of 1.95 [75], while it decreases when fracturing alters the flow front into a fingering-like pattern (B3). Similarly, the fractal dimension of the fracturing finger decreases as increases, consistent with the experimental range [1.70, 1.82] for enhanced fracture propagation with increasing [75]. Overall, our model effectively captures the transition from an infiltration-dominated regime to a fracturing-dominated regime as increases, providing a quantitative representation of the evolution of the ramified flow front and fracturing channels.
@@ 5 Conclusions
Moreover, our model offers quantitative insights into the underlying micro-mechanisms governing multiphase flow in porous media, marking the first time that micro-scale interactions have been visually reviewed in terms of viscous and capillary pressures, energies, contact forces, and flow resistances. In simulating Lenormand’s experiment, we identify pattern-transition signals from viscous fingering and stable displacement to capillary fingering by comparing throat-scale viscous and capillary entry pressures. In simulating Huang’s experiment, we elucidate the governing regimes of fluid–grain interactions by comparing pore-scale residual energies and contact forces for fracturing, and flow resistances for fluid dissipation.
Notably, while high flow rates and increased invading fluid viscosity in deformable porous media enhance fracturing and influence fluid infiltration paths, this effect contrasts with the positive impact on production efficiency observed in rigid porous media. This discrepancy highlights the need for further systematic parametric studies using our model to fully understand the trade-offs between operational efficiency and system stability in geological processes. Future work will explore the role of gas compressibility in optimizing gas storage and recovery, leveraging our model’s unique capability to account for fluid compressibility—a feature not addressed by most existing models, which often assume incompressible or overly simplified fluids. Additionally, we aim to enhance the model by replacing fixed boundaries with confined ones, enabling more realistic simulations of multiphase flow patterns and displacement efficiency under stress-dependent conditions.
@@ Acknowledgments
This work was supported by the HKU Postgraduate Scholarship. Kand Duan acknowledges funding from the National Key RD Program of China (No. 2023YFB2390300). We thank Dr. Yue (Olivia) Meng (Purdue University) and Dr. Zhibing Yang (Wuhan University) for insightful discussions that contributed to this study.
@@
Appendix A Resolution convergence
We here describe additional tests conducted to ensure that our validation samples have sufficient resolution, with enough particles to produce consistent results. The key criterion for resolution convergence is that flow patterns and temporal pressure evolution remain consistent across different resolutions, achieved by varying grain sizes while maintaining constant permeability and operational parameters ( and ). This ensures that viscous pressure and capillary entry pressure balance at similar levels within the displacement regime.
We use Lenormand’s fixed-grain case, where air displaces a very viscous oil (Case 1), as an example. Specifically, we vary the original average grain radius () by decreasing it by 30 ( mm) and increasing it by 30 ( mm), resulting in particle counts of 2305 (refined) and 724 (coarser), respectively. The sample size, random control interval (), and boundary conditions are kept the same as in Figure 4. The corresponding refined and coarser pore structures are shown in Figure 17. The original pipe network, consisting of 1874 domains and 3061 pipes, is refined to 3403 domains and 5707 pipes for the smaller grain size and coarsened to 1160 domains and 1883 pipes for the larger grain size. To ensure that permeability remains consistent across all samples, the aperture width range is adjusted proportionally. For the refined sample, the average aperture width decreases to 0.35 mm, while for the coarser sample, it increases to 0.47 mm, compared to 0.42 mm in the original case. These adjustments ensure that all samples achieve the experimental permeability of 10-9 m2.
As shown in Figure 17, the simulated flow patterns exhibit key features of viscous fingering (tree-like morphology) and capillary fingering (thicker fingers with blocked front pipes) for both the reduced- and increased-grain-size samples. These observations are consistent with the original mm case, as shown in Figures 5a and 5b. We also compare the injection pressure development between the original sample and the grain-size-modified samples for the viscous fingering regime. As illustrated in Figure 18, the pressure curves exhibit excellent agreement in both magnitude and temporal trends, confirming resolution convergence and demonstrating the robustness of our simulation framework across variations in pore structure.
Appendix B Timestep sensitivity
We here investigate the sensitivity of our numerical validations of Lenormand’s and Huang’s experiments to timestep variations. Specifically, we take viscous fingering (VF: ), capillary fingering (CF: ), fluid infiltration-dominated stable displacement (B1: ), and fracturing-dominated stable displacement (B3: ) as representative examples. For these cases, we vary the timesteps by increasing or decreasing them by 50. For instance, in CF of Lenormand’s experiment, where the original timestep is 100 s, additional simulations are conducted with s and s. The model setups remain consistent with those shown in Figure 4 for fixed-grain injection cases (Lenormand’s experiment) and Figure 11 for movable-grain injection cases (Huang’s experiment).
As shown in Figures 19 and 20, both the fluid-fluid displacement patterns (i.e. viscous fingering and capillary fingering) and fluid-grain displacement patterns (i.e. fluid infiltration-dominated and fracturing-dominated regimes) closely match the original validation results of Lenormand’s (Figure 5) and Huang’s experiments (Figure 13). The temporal evolution of injection pressure also agrees well between the additional tests and the original simulations, in terms of pressure magnitude and development trends. The only notable variation occurs in the peak injection pressure of B3, which features a high inflow rate and invading fluid viscosity. When the timestep is increased from to s, the peak injection pressure changes from 9.0 to 11.1 MPa. This variation is attributed to differences in the frequency of updating the mixing water-air fluid bulk modulus in the injection domain under varying timesteps. While this change may influence the early stages of fracture initiation and propagation, it does not impact the final fluid-grain displacement pattern.
The use of larger timesteps in our fully coupled HM-DEM framework is inherently not problematic. Our default flow front-advancing criteria (Eqs. (18) and (19)) automatically refine the timestep to ensure stability when any front domain becomes overfilled or when flow is impeded due to blocked front pipes caused by slow drainage. However, excessively small timesteps reduce simulation efficiency, particularly for fluid infiltration-dominated stable displacement (B1), as shown in Table 5. In summary, all our validation results demonstrate both accuracy and robustness with respect to timestep variations.
| Lenormand-VF | s | s | s |
| [hour] | 1.2 | 1.0 | 1.0 |
| Lenormand-CF | s | s | s |
| [hour] | 7.0 | 7.5 | 8.3 |
| Huang-B1 | s | s | s |
| [hour] | 45.5 | 16.2 | 7.2 |
| Huang-B3 | s | s | s |
| [hour] | 7.5 | 2.0 | 1.5 |
Appendix C Pressure solver performance
In this study, we apply an implicit pressure solver to overcome the timestep limitations inherent in explicit integration schemes commonly used in pore network models (e.g. [79, 17]). To demonstrate its performance improvement, we conduct additional simulations using an explicit pressure solver and compare the results with those obtained using our implicit solver. Representative cases from Lenormand’s and Huang’s experiments—including viscous fingering (VF: ), capillary fingering (CF: ), fluid infiltration-dominated stable displacement (B1: ), and fracturing-dominated stable displacement (B3: )—are used for this comparison.
Using an explicit time integration scheme to solve the fluid mass balance (Eq. (2)) necessitates a very small timestep for stability [17]. Ignoring domain volume changes, the characteristic timestep for fluid dynamics within a domain is given by
where is the pressure variation per timestep. The explicit form of the Plane Poiseuille flow equation (Eq. (1)) is approximated as
Substituting this expression into the fluid mass balance and applying the scalings , , and (with the average grain radius), the critical timestep becomes
where is the number of pipes per domain (assumed to be 5 to represent the critical situation). Given the extremely small critical timestep, the mechanical timestep in explicit simulations is set equal to the fluid flow timestep.
Table 6 summarizes the performance comparison between implicit and explicit pressure solvers, including fluid flow timestep (), flow-to-mechanical timestep ratio (), total simulation time (), computational time (), number of run steps, and solver efficiency (dimensionless time: ). The number of steps also accounts for refined timesteps required to manage overfilled domains and blocked front pipes. Although the explicit solver executes more steps within the same computational time—largely due to fewer mechanical steps and the reduced computational load of solving the pressure matrix—this apparent efficiency is misleading. Closer examination reveals that the dimensionless time for the explicit solver is orders of magnitude smaller than that of the implicit solver. Specifically, the implicit solver outperforms the explicit solver by roughly nine orders of magnitude in CF () and two orders of magnitude in B3 (). This disparity arises because the explicit solver’s extremely small timesteps severely limit the progress of fluid dynamics per step, ultimately rendering it computationally inefficient.
Flow pattern comparisons (Figure 21) further emphasize this point. Despite executing far more steps, the explicit pressure solver barely advances the invading fluid within the injection domain. In contrast, the implicit solver, with far fewer steps, allows the invading fluid to reach the outflow boundary within the same computational time. This demonstrates the implicit solver’s ability to utilize much larger timesteps, offering a clear computational advantage.
In summary, while the explicit solver executes more steps, the implicit solver’s ability to handle larger timesteps and generate realistic multiphase flow patterns makes it a significantly more efficient and practical choice for simulating multiphase flow in deformable porous media.
| Test | [s] | [s] | [hour] | Steps | / | |
| VF-implicit | [-] | 10.5 | 1 | 703 | ||
| VF-explicit | [-] | 1 | 4010 | |||
| CF-implicit | [-] | 9324.1 | 7.5 | 1783 | ||
| CF-explicit | [-] | 7.5 | 27555 | |||
| B1-implicit | 16.2 | 2910 | ||||
| B1-explicit | 16.2 | 12415 | ||||
| B3-implicit | 2 | 612 | ||||
| B3-explicit | 2 | 2211 |
Appendix D Simple flow tests
While our validation results using randomly packed samples align well with experimental observations in both Lenormand’s and Huang’s experiments (Figures 5 and 13), we further evaluate the model’s accuracy through two simple flow tests: single-phase Darcy flow in randomly packed media and two-phase flow in regularly packed media. Unlike multiphase fluid injection in highly randomized samples, Darcy flow can be validated against analytical solutions, and two-phase flow in regular packing allows precise reproduction of experimental results, such as the number and direction of fingers.
D.1 Single-phase Darcy flow test
We simulate one-dimensional (1D) single-phase linear flow by injecting highly viscous oil (same as Case 1 in Table 1) into oil-saturated, fixed-grain porous media. The model setup is the same as in Lenormand’s experiment (Figure 4), with an injection rate of m3/s applied at the left boundary. The simulation continues until inflow and outflow reach steady state.
The simulated results are validated against analytical solutions for 1D Darcy-flow problem, governed by the diffusion equation
with the dimensionless quantities for pressure , -coordinate distance from the inflow boundary , and time defined, respectively, by
Here, is the intrinsic permeability, is the initial pressure, is the consolidation coefficient, is fluid viscosity, and are the height and length of the porous media, is the out-of-plane thickness (1 m for 2D problems), and is the total inflow rate.
The initial and boundary conditions are
and
We derive analytical solutions for fluid pressure and flow rate by the method of separation [66], which can be expressed as
and
Theoretical inflow and outflow rates are evaluated at and , respectively. Permeability is determined from the steady-state pressure gradient and flow rate using Eq. (21), while is calibrated by matching simulated inflow and outflow curves at steady state.
Figure 22 compares simulation results with analytical solutions. The simulated inflow and outflow rates match closely with analytical solutions when (Figure 22a), confirming the consistency between theory and simulation during transient and steady-state conditions. Similarly, the simulated pressure profiles align with analytical predictions at all selected time stages (Figure 22b). Despite minor randomness introduced by variations in domain and pipe sizes, the results demonstrate excellent agreement.
These findings validate the model’s ability to accurately simulate fluid migration in permeable media, confirming its reliability and robustness in Darcy-flow analysis.
D.2 Two-phase flow in regular packing
In random packing, two-phase flow patterns are typically classified based on morphological features (e.g. tree-like structures or blocked front pipes) or fractal dimensions. In contrast, regular packing, while overly simple for representing real geological formations, allows flow patterns such as the number and direction of fingers to be reproduced exactly as observed in experiments.
To validate this, we replicate a viscous fingering experiment conducted by Chen and Wilkinson [80], where non-wetting oil ( cP) is injected into wetting glycerine ( cP) within a square network of equally spaced, interconnected channels etched into glass. While the original experiment did not measure permeability, we set the grain radius () and aperture width () to 0.25 mm, matching experimental values.
Our model constructs a 39 39 square network of channels, each 0.5 mm in length, consisting of 1521 domains and 3120 pipes (Figure 23). The inflow domain is located at the center, while 72 outflow domains along the edges are maintained at a constant zero pressure. All initial domain pressures are set to zero. The fluid bulk modulus is 2000 MPa for both oil and glycerine, with interfacial tension of 20 dyn/cm and a contact angle of 180.
We simulate drainage with a modified capillary number , corresponding to the experimental inflow rate of m3/s, replicating the experiment’s viscous fingering behavior. As shown in Figure 23, the simulation reproduces the observed pattern, forming four fingers along the -axes, identical in number and direction to the experiment.
This accuracy arises from the uniformity of pore and aperture sizes, which ensures consistent pressure distribution favoring flow along the -axes, resulting in the formation of four fingers. This successful reproduction underscores the accuracy of our HM-DEM framework in simulating two-phase flow in regular media, providing a reliable benchmark for validating its application to more complex systems such as Lenormand’s and Huang’s experiments.
Appendix E Fractal dimension calculation
The classification of various fluid-fluid and fluid-grain displacement patterns is often based on visual inspection of their characteristic morphological features, such as the highly branched viscous fingering or the less-branched capillary fingering. While such qualitative approach can be useful, we also adopt a quantitative method to classify displacement patterns based on their fractal dimension, , calculated via the box-counting method [71]—a widely used technique for quantifying the complexity of irregular shapes or flow patterns. In this method, the fractal domain is discretized into a grid of square boxes with side length , and the minimum number of boxes, , required to cover the domain is counted. The process is repeated for a range of box sizes, and the fractal dimension is determined as the slope of the linear region in a log–log plot of versus :
In practice, the resolution of the image or the pore structure in a pipe network model (e.g. grain and pore sizes) can greatly influence the accuracy of fractal dimension estimation. To address this issue, we employ a “local fractal dimension” approach. Instead of fitting a single slope across all scales, we calculate the slope over smaller intervals in the log-log space and then average these local slopes over the scales most relevant to the observed patterns. This approach enhances robustness by accounting for variations in resolution and focusing on the box sizes that best capture the dominant morphological features, providing a more reliable estimate of the fractal dimension for fluid-fluid and fluid-grain displacement patterns.
Before applying the box-counting method, the flow pattern image is first binarized to isolate the infiltration or fracturing regions. Subsequently, we plot versus as well as the local versus to determine the fractal dimension of the specific flow pattern while accounting for pore structure resolution. As a demonstration, we calculate the fractal dimension for three representative cases: viscous fingering () and capillary fingering () from Lenormand’s experiment, and the fracturing-dominated stable displacement case (B3) from Huang’s experiment. The results are shown in Figures 24 and 25. A summary of the calculated fractal dimensions for all simulated displacement patterns is provided in Figure 26.